OpenStudy (anonymous):

If you have 240.0 mL of water at 25.00 °C and add 100.0 mL of water at 95.00 °C, what is the final temperature of the mixture? Use 1.00 g/mL as the density of water.

OpenStudy (adamconner):

heat gained + heat lost =0 Q = mass x specific heat x delta T 240mL x 1 g/mL = 240 g 100 mL x 1 g/mL = 100 g 240 x 4.18 ( T - 25) + 100 x 4.18 ( T - 95)=0 divide by 4.18 ( specific heat of water) 240 ( T - 25 ) + 100 ( T - 95)=0 240T-(240x25) + 100T-(100x25)=0 240T -6000 + 100T- 2500=0 240T+100T=6000+2500 340T=8500 T=8500/340 T=25C

OpenStudy (anonymous):

Thank you so much!

OpenStudy (adamconner):

your welcome!!